Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability that the marble selected will be purple?

Expert replies
by rsarashi » Sat Sep 02, 2017 11:49 am
On a store counter are exactly two boxes containing only purple marbles and yellow marbles. Box A has 30 purple marbles and 20 yellow marbles; box B has 15 purple marbles and 35 yellow marbles. If Melanie randomly selects one of the boxes and then randomly selects one marble from that box, what is the probability that the marble selected will be purple?

A) 3/20

B) 3/10

C) 9/20

D) 3/5

E) 9/10

OAC
Join the discussion
Source: — Problem Solving |

by Brent@GMATPrepNow » Sat Sep 02, 2017 11:59 am
rsarashi wrote:On a store counter are exactly two boxes containing only purple marbles and yellow marbles. Box A has 30 purple marbles and 20 yellow marbles; box B has 15 purple marbles and 35 yellow marbles. If Melanie randomly selects one of the boxes and then randomly selects one marble from that box, what is the probability that the marble selected will be purple?

A) 3/20
B) 3/10
C) 9/20
D) 3/5
E) 9/10
P(marble is purple) = P(Box A is selected AND a purple ball is selected from Box A OR Box B is selected AND a purple ball is selected from Box B)
= P(Box A is selected) x P(a purple ball is selected from Box A) + P(Box B is selected) x P(a purple ball is selected from Box B)
= 1/2 x 30/50 + 1/2 x 15/50
= 3/10 + 3/20
= 6/20 + 3/20
= 9/20

Answer: C
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by GMATGuruNY » Sat Sep 02, 2017 1:25 pm
On a store counter are exactly two boxes containing only purple marbles and yellow marbles.Box A has 30 purple marbles and 20 yellow marbles;box B has 15 purple marbles and 35 yellow marbles. If Melanie randomly selects one marble from any of the boxes ,what is the probability that the marble selected will be purple?

A)3/20
B)3/10
C)9/20
D)3/5
E)9/10
The two boxes are irrelevant.
Total number of purple marbles = 30+15 = 45.
Total number of marbles = 30+20+15+35 = 100.
P(purple is selected) = purple/total = 45/100 = 9/20.

The correct answer is C.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

rsarashi wrote: ↑
Sat Sep 02, 2017 11:49 am
On a store counter are exactly two boxes containing only purple marbles and yellow marbles. Box A has 30 purple marbles and 20 yellow marbles; box B has 15 purple marbles and 35 yellow marbles. If Melanie randomly selects one of the boxes and then randomly selects one marble from that box, what is the probability that the marble selected will be purple?

A) 3/20

B) 3/10

C) 9/20

D) 3/5

E) 9/10

OAC
We are given that box A has 30 purple marbles and 20 yellow marbles and that box B has 15 purple marbles and 35 yellow marbles. We need to determine the probability, when selecting one marble from the box, of selecting a purple marble. Since we have two boxes, A and B, there are multiple scenarios to account for when selecting the marbles. We must account for the probability of first selecting either box and then secondly for selecting a purple marble. Let’s start with Box A.

P(selecting box A) = ½

P(selecting a purple marble in box A) = 30/50 = ⅗

Thus, the probability of selecting a purple marble from box A is ½ x ⅗ = 3/10

Next we can determine the probability of selecting a purple marble from box B.

P(selecting box B) = ½

P(selecting a purple marble in box B) = 15/50 = 3/10

Thus, the probability of selecting a purple marble from box B is ½ x 3/10 = 3/20

Now we can determine the probability of selecting a purple marble from box A or box B:

3/10 + 3/20 = 6/20 + 3/20 = 9/20.

Answer: C

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion